Country Foreign currency units per U.S. dollar Year 1 Year 2 Brazil (real) 2.379 2.088 China...

Country

Foreign currency units per U.S. dollar

Year 1

Year 2

Brazil (real)

2.379

2.088

China (yuan)

8.277

8.004

Mexico (peso)

9.088

11.393

Venezuela (bolivares)

717.27

2144.60

You and a friend, Ying, are considering a summer vacation to one of the two locales: Mexico or Brazil. Mexico’s currency unit is the peso and Brazil’s currency is the real. You live in the United Sates and Ying lives in China. The cost of the hotel room is denominated in local currency units. You and Ying are trying to find the locale with the lowest hotel expenses. The price of a hotel room is 200 reals per night in Brazil and 800 pesos per night in Mexico. Assume these prices remain unchanged between Year 1 and Year 2.

  1. Calculate the U.S. dollar price of a hotel room in Brazil and in Mexico for Year 1. Calculate these prices for Year 2. Based on your answers, where would you prefer to spend your vacation in Year 1? And, where would you prefer to spend your vacation in Year 2?
  2. For Year 1 and Year 2, calculate the Real-Yuan and the Peso-Yuan exchange rates.
  3. Using your answer from (b) above and the information in the table, calculate the Chinese yuan price of a hotel room in Brazil and in Mexico for Year 1. Calculate these prices for Year 2. Based on your answers, where would Ying prefer to spend her vacation in Year 1, and in Year 2?
Answer #1

a) I live in the U.S.

In year 1,

given $1=2.379 real.

and $1=9.088 peso.

Therefore, In year 1 if i have to spend a night in Brazil,I would have to pay $(200/2.379)=$84.069

and in year 1, If I have to spend the night in Mexico, would pay $800/9.088=$88.028

Therefore in year 1,I would prefer to spend the night in Brazil.

In year 2 ,

given $i=2.088 real.

and $1=11.393 peso.

Therefore, In year 2 if i have to spend a night in Brazil,I would have to pay $(200/2.088)=$95.785

and in year 2, If I have to spend the night in Mexico, would pay $800/11.393=$70.218

Therefore in year 2,I would prefer to spend the night in Mexico.

b)For year 1----

$1=8.277 yuan and $1=2.379 real

or 8.277 yuan=2.379 real

or 1yuan=2.379/8.277 real=.2874 real

real/yuan=.2874

$1=9.088 peso and $1=8.277 yuan

or 8.277 yuan=9.088 peso

or 1 yuan=9.088/8.277 peso=1.098 peso.

Therefore,peso/yuan=1.098

For year 2----

$1=8.004 yuan and $1=2.088 real

or 8.004 yuan=2.088 real

or 1yuan=2.088/8.004 real=.2609 real

real/yuan=.2609

$1=11.393 peso and $1=8.044 yuan

or 8.044 yuan=11.393 peso

or 1 yuan=11.393/8.044 peso=1.416 peso.

Therefore,peso/yuan=1.416

c)Price of a hotel in Brazil=200 real

Price of hotel in Mexico=800peso.

In year 1,

amount of yuan required to spend a night in Brazil=200/.2874=695.894 yuan

amount of yuan required to spend a night in Mexico=800/1.098=728.597 yuan

Therefore,Ying prefers a night in Brazil.

In year 2,

amount of yuan required to spend a night in Brazil=200/.2609= 766.577 yuan

amount of yuan required to spend a night in Mexico=800/1.416= 564.972 yuan

In year 2, Ying prefers a night in Mexico.